NAOJ GW Elog Logbook 3.2

Yesterday I switched on the optical lever of the end mirror. The signals in time were showing a clear excess of noise with respect to the last time I have looked at them (july 2016).
Since the end chamber was put under vacuum few weeks ago, we suspeced the pumps. We first switched of the turbo pump (after closing the valve) and wait for the frequency to go to zero without notice any remarkable improvement in the noise. After that, we switched off the rotary pump and we could observe that many lines have disapeared in the spectrum. Indeed the signal in time was much better than before.
We also monitored the change in the vacuum level after closing the valve. Its trend in the first 45 minutes is plotted in the attached figure.
The value today (after about 18 h) is 2.1e-4 torr.

This morning I have dismounted the gluening jigs that we used to glue magnets and standoff on the end mirror 2 days ago. Magnets and standoff seem well attached but while trying to put the mirror inside its case a magnet came off. We reglued it just after cleaning the mirror. After few hours we removed the jig and let the glue cure without support (picure1)
We then proceeded measuring the wedge (pictures 2-3) and glueing the magnets on the input mirror (support number 4), following the same procedure reported here.
Participants: Eleonora, Marc
Few days ago we observed that on one of the stand-off we glued on the input mirror the glue seemed to have overfowded on the groove where the wire is supposed to stay (picture1). Today we made some tests with a wire of the same diameter of those used in the suspension and verified that it was the case. After trying to remove the excess of glue without good results, we decided that was safer to remove the stand-off and to glue a new one. (picture 2)

There are a total of six pressure measurements along the south arm.
- Near NM2
P1 = None (NM2 is not evacuated)
P2 = 2.5e-8 Torr
- Mid-arm
P1 = 2.0e-8 Torr
P2 = 2.5e-8 Torr
- Near EM2
P1 = 1.2e-7 Torr
P2 = 1.0e-7 Torr

Participants: Eleonora, Marc, Raffaele
Filter cavity mirrors are supposed to have a wedge of 400 urad. (See picture 1)
Since the wedge is not marked on the edge of the mirrors (as it happens usually), we had to measured it using an autocollimator. This tool works by projecting an image (a cross) onto the mirror and measuring the deflection of the returned image against a screen with a grid. If the mirror has a wedge, the reflection of the first and second surface are not superposed, resulting in two crosses on the screen. The line joining the crosses' center indicates the wedge direction (i.e the diameter with the maximum slope). It is not easy to precisely project the wedge line seen on the screen, on the mirror surface in order to glue the stand offs in the proper position (we want the wedge to be horizontal). In order to do this we used the following procedure:
1) We used as a reference a 4" spare mirror (used so far for gluing test) which as a marked wedge (figure 2)
2) We put it on the glueing support and aligned the wedge with the stand off jig.
3) We moved the autocollimator in order to have the two crosses well aligned on an axis of the screen grid.
4) We replace the dummy mirror with the filter cavity mirror and rotate it in order to have the crossed aligned on the same axis as before. In this way the wedge should be aligned with the stand off glueing jig.
Since we can not distiguish the first and the second reflection of the mirror we don't know if the wedge is such that an impinging beam is reflected on the right or on the left.
We proceded glueing the magnets and stand-off. We glued the end mirror number 1 that, according to my simulation (see figure 4), is the best one to be used together with input mirror number 4.
We used a brand-new set of masterbond glue and we did all the work in the ATC clean room.
-------------------------------------------------------------------------------------------------------------------------------------------
We have measured a distance between the crossed of almost 2 division of the screen grid. We are not still sure 100% about how this data should be convert into a wedge value. Marc will report on it as soon as we find it out.

Based on this result and the aperture of the optics, we design the telescope:
Origin set at the first line of holes in front of the cavity.
L1: z=7.5cm f1=100mm
L2ï¼›z=45cm f2=125mm
L3: z=132.5cm f3=1000mm
According to the optical scheme, we estimated the position of the optics and their apertures:
FI: 22.5cm 5mm
EOM:57.5cm 3mm
AOM:90cm 1.5mm
FI: 123.5 5mm
In the simulation, after we put the three lenses for the telescope, we only need to put another two tune lenses for the AOM, two 100mm, one at 82.5cm, the other at 102.5cm. Then everything will be fine.
Another thing we discovered today, it's that the telescope for infrared has been finished for very long time, but with a lens having 350mm focal length which is not available from any company, so we did the design again:
L1: z=48.5cm f1=500mm(Alreday installed)
L2: z=122.5cm f2=150mm
L3: z=154.5cm f3=175cm
We will try to look into the bench if there is enough space for the lens and order them.
Also yesterday we put everything back one by one and send the beam far away to check where exactly the astigmatism come from. Everything was find until we put the Faraday, then after the mirror we put the second lens, when we moved this lens or turned it, the beam can change from elongated in vertical to round then from round to elongated in horizontal. So finally we can be sure that the astigmatism is from this lens. Then we put it in a good position and tried to send it to the end. On the screen, we can see it on the 290m target, although it is large, but it more circular than before, we will check the size of it after we finish the telescope,if it is still large, we can move the 2inch mirror.

At the beginning, we found out that the beam shape in the beam profiler moved a lot, it was a ellipse with a very large e. But if we sent the beam far away from the bench, it was round,(pic 1), but only with some fringe on one side, it seems from the dichroic mirror.
Then we tried to turn the dichroic mirror a little to see if the beam change. Since the dichroic is also working for sending the infrared beam into the cavity, we moved it, so we need to align the infrared beam to get the green back. Then we also find out the green power is low, so we adjusted the thermal control to find a good temperature, finally we got 50mW over 240mW of infrared, efficiency around 20%.
When everything was ready, we found out the beam size is much smaller compared to the data we got last week, so maybe the temperature also effect the beam size of the green, we took several points, did the calculation, this time the beam waist has more or less the same position of the infrared(-7.3cm),but half of the size(26 micrometer). In picture two it is the points we got in two axis of the beam profiler, and picture three it is the data points and the fitting result in one of the axis.
Also we check again the polarization just after the cavity, we got 49mW reflected by the PBS and 1mW transmitted, so we can say that the green beam come out from the cavity is in S polarization which we want it be.
We did these measurement with the air conditioner off, since the wind from all the air conditioner meet at the bench, produced a lot of fluctuation in the green beam. Then when we left, we turned the air conditioner on,which means we need to adjust the thermal control again.
The other thing is after we realigned the cavity, we found every time we locked the cavity, there is the high frequency oscillation, so we reduced the gain of the Stanford from 500 to 50, then we are fine now.

Members: Manuel, Yuefan, Marc
- We made order in Tama central room, sorted many things from the messy plastic boxes and arranged them on the new shelf.
- We clean up some dust from the cleanbooth of the absorption bench and move the translation stage on the definitive optical table, see picture.

So today we remove the two dichroic, the lens and Faraday, finally we had enough space to put the beam profiler and do some measurement, but even like this, we only can took three points,after the third point, we still had space but the beam size is too large to measure. The beam profiler has two axis,v and w, the first time we took three data points, two of them is fine, the beam size in two axis is similar, but one point had a large difference. So after the calculation, the beam waist and position in these two axis is very difference(one has half size of the other). Then we did the second measurement, this time the result is good, and we got the beam waist size was 36 micrometer instead of 52 micrometer(the beam waist size of the infrared), and the beam position is 15.5cm away from the input port of cavity instead of 7.5cm(infrared).
Although we got this result, we still have the suspect that maybe the astigmatism is from the cavity, since the beam size is too large, the measurement may have some error.
Then because we already remove the dichroic, we tried to turn it to make the two beams from two surface of the mirror can be separated. But we turned the mirror little by little almost 360 degree, the two beam still become into one. So we thought maybe the wedge is too small, even it separate the two beam, but the beam size increase too quickly, then the main beam will cover the other.
Before the holiday, we found out the picomotor which controls the yaw of BS mirror cannot be controlled by the computer. And we checked the cable, from the driver to the connected part of the picomotor and the cable, it works well. Also we exchanged the port between pitch and yaw, after pitch works well, so the problem is also not from the driver. So Marc checked again the picomotor, he thought the problem is inside the picomotor,and we planned to change the it after we finished the green path and need to align the beam.

But if we change to one with 50mm focal length,there is a possibility that the before the beam goes out from the Faraday it will expand. So I did the simulation today. I used the initial beam waist is 52.7 micrometer and 82.2cm away from the output of EOM, this is the value we had from the infrared beam, I am not sure if the green beam has the same size and position as the infrared one, but if I did the simulation with 100mm lens in its position now, the simulation result is close to the real beam size on the bench, so I did the after simulation with this starting point. Then I tried to change the lens to 50mm one in the simulation, and got the result in the first picture. The length of the Faraday is about 5cm, so if I put the beam waist at the center of the Faraday, the beam size at the input and output port of the it will be 762 micrometer in diameter which I think it is small enough.
But if we changed like this, there may be another problem, the aberration on the first lens will be much serious, since the beam size on the first lens does not change, but the focal length reduce.
We got the PBS for green today, so we can check the polarization of the green beam. If I remember right, we got pure P polarization went into the cavity, so the green should be S polarization, so we adjust the waveplate to maximum the S polarization.

Yesterday afternoon we tried to lower the inter-medium mass first, but we round out one of the top four screw is very hard to loose(pic 1), so we tried to lower the mass with one of the picomotor. But only with the picomotor the wire was not loose enough to remove the mirror, so finally we managed to turn that screw, and today when we need to fasten it, we changed to another screw.
So today we tried to put the real mirror on the wire, the stand-off and the magnet are all stick well, but since the real mirror is a little bit smaller, the white stopper on the up side of the mirror can not help a lot. Also when we try to put some intense on the wire to lift if from the stoppers, the mirror tilt a lot, so we turn the pitch picomotor to make it more or less vertical.
Then we tried with picomotor to align the beam again, while we were trying we found out the yaw picomotor of BS suspension does not work, we checked a bit, for now we can only say it is not the problem of the driver.We will try to fix this problem tomorrow morning. Then we tried to align with move the picomotor by hand, so we did align the beam, but maybe because the quality of mirror becomes higher, there are too many reflection everywhere, so the beam on the 290m target looks much worse than before.(pic 2)
The other thing is that we checked the picomotor driver for the end room yesterday and also checked again today, it works well both the top layer and the port layer. Now we have enough drivers.

Yesterday we move the optical bench in the NM2 chamber in order to have the beam
approximately passing on the beach axis. To do so the beam was previously positioned so to
pass approximately at the center of the two windows on the gate valves in the center area
(between the BS chamber and the NM2 chamber and between the NM2 chamber and the tube)
and reaching the target at 290 m.

With putting two aperture on the bench (pic 1), we set the position of the beam, and started to move the lens, at first when we just turn the screw in horizontal direction to center the beam on the lens, the beam is still elongated. Then we tried to turn the lens a little bit,then we found out the elongated changed to the vertical direction. If look at picture 2, it is not very obvious, but when we tried to send the beam far away, the beam is longer in the vertical direction a lot. We thought the reason of the beam shape is because of the large beam and short focal length(150mm), since the spherical aberration, when the beam pass the spherical surface, the light away from the axis does not focus on the same point as the light along the axis, so the more we off the center of the lens, the beam elongated more. So then we will try to center the beam better on the lens.
But the other strange thing is that the beam at 290m target is still elongated on horizontal direction(pic 3), this situation is the same when we tried to use the aperture to force the beam to beam round, then at the first target it looks good, but then on the second target it is pretty long in horizontal. We need to have a better beam and then try to find the possible reason of this.
The other things we discovered on Friday is that why when we center on all the mirror and also second target, we were off center of the first target. Then first target has a axial precession like the Earth's axis, so if we turn it to the other face it will be well centered also.

Then after we put the mirror like the type for the local control,working well in vacuum, to reflect the beams come out all from one window, on the right side of the input window of PR if you face the optical bench. Then we use a plastic sheet to put some cross on the position of the beam,and fixed them near the window( Pic 2), we use this kind of mount then if we need we can remove the plastic and then put them back again on the right position.
Yesterday we suddenly lost the infrared alignment, the beam went into the chamber but missed the first mirror. We checked the beam on the bench, there are totally three mirror to turning the beam into the chamber. Between the 1st and 2nd, we check with the target, it is fine. Then we wanted to know the reason why suddenly we lost the alignment, so we try to push a little bit at the edge of the 2nd and 3rd mirror, it seems the 2nd mirror fixed very good on the mount, but the third one does not. Also the beam is near the left edge of the 3rd mirror, and also we missed the 1st mirror in the chamber with the beam on the left side. If we wanted to move the beam on to the 1st mirror in the chamber, we need go more left on the 3rd mirror, then we will go out of this one. So we decided to move only the third mirror. Luckily, we got the alignment back with only moving this one.
Also we put another mirror after the dichroic mirror to check the reflect infrared beam can go to the right point on the bench or not. So it seems the alignment is very good, the infrared go back through the Faraday and reach good position.
We finished everything of the PR chamber, so in the afternoon we moved the clean booth to the BS chamber and closed the PR.(pic 3)
Today we also work a little bit on the astigmatism. At first we only want to put some aperture to set the beam position on the bench for re-align easier in the future. Then we found out if we make the beam on the bench through the aperture to make them round (pic 4), the beam after the telescope looks pretty good.(pic 5). So this time we were sure that the astigmatism is not from the things in chamber. Then we tried to put another aperture after the lens on the bench, the large beam looks very good as well, then put it before the lens, it seems the beam got worse, but since the beam before the lens is very small, it was hard to say the lens is the only reason for that.
Pictures of the opened PR chamber.
3 mirrors have been added to obtain the IR and green references.
G1 is the green beam reflected by the dichroic
IR 1 is the IR beam after the Faraday Isolator and a folding mirror
GR2 and IR2 are reflected by the same mirror ( between the green folding mirror and the dichroic and between the 2 metallical structures).
They can be seen on the "PR references" picture.
I will try to upload a more precise optical scheme of this chamber

Also the SHG cavity had high frequency oscillation when it went around the resonance, so we changed the low pass filter frequency to 3Hz and the gain to 200.
The other problem we found out yesterday is that the optimal temperature of the SHG has been changed, before the thermal controller was set with the resistance of 3.270 where we can get the maximum power of green. But yesterday we found out the green beam power is too low, but the mode matching seems not bad, so we tried to change the temperature, and now the maximum green power was at 3.230, there we have the power about 20mW, it is still lower than before. One of the possibility is that the environment temperature in TAMA increased recently, so the radiator may not work as good as before.
And also now TAMA air conditioner is working at winter mode, when we tried to change it to summer mode, we found out the green beam shaked a lot. It may caused by the air flow between different temperature, since TAMA is quiet large, maybe it will take a long time for the temperature to get stable, so we planned to set the air conditioner to summer mode after we finish work this Friday, let it adjusts along all the weekend, hope on Monday it is stable enough for us to work on.

Then we went to the end room, the beam has already came out from the EM chamber, we saw it on the wall. After that we took a camera to the end room to look at the beam on the wall, connected to the board there and received the signal in the central room(pic2). Since the beam is shaking a lot because the air flow in the central room, so the picture is not very clear.
Although we sent the beam to the end mirror, but the beam seems larger then the one we sent. We are going to move the 2 inch mirror in the BS chamber tomorrow, try to find a good focus position.

So from the first picture you can see the blue pipe, it is connected to a pump, then it can move the platform in picture two to move the target up and down. Then in the second picture, there is a motor used to turn the target. So both of these two controller are connected to the blue box in picture three. From what written on this box, we found out we can control this system locally and also remotely. So we followed the cable from this box to the end room, found another controller there(Pic 4&5). Now we keep the target there with moving the platform and motor locally, but if we want to make future work easier, it is better to find out how to do the remote control with the controller in the end room.

First in the situation as last Friday, we checked the reflected beam from the window, it was almost like before, very obvious elongated in horizontal direction, and also checked after the window, it seems even worse than before, it used to be about circular.
Then we tried to turn the dichroic mirror in the PR chamber to reflect the infrared beam enough far(into the west arm,PI) to look at the shape, it is more or less round beam.
The third beam is the transmitted of BS1, we checked across the table(P1), it is elliptical.
Next one we checked the transmitted beam of M2, luckily this beam pass successfully through the about 5cm gap between two tubes.We had enough distance to check it(P2), this beam is already started to be longer in horizontal.
Then the last check we put another mirror between L1 and M2(MT) to reflect the beam,checked with almost the same distance with P1, it is not a perfect round beam from my sight. But then we put the beam profiler at the edge of the bench, try to get the shape with it. Actually from the profiler, it seems a pretty good beam.
Then since we already finished all the picomotors, and I finished to make the cover of the extended wire today, so then we tested the extended cable again in case that I broke some during making the cover. All of them works well. But when we tried to move the picomotors of PR chamber, there is one layer of dirvers did not work. So now we have one problem is that we do not have enough drivers for the end room.
After fixed all the problems of picomotor, then we tried to re-align the beam. We recovered the last mirror on the bench for green first(M4), since we set the position with the target, it was not very hard to do it. But then we found out that the green beam on both the turning mirror and the dichroic mirror are both not very centered, I am not sure about the situation after the middle part of the chamber put back, but since we did not change M4 in horizontal, so we keep working with that. The we realign all the beam until the BS suspension mirror.
We took the 2cm shift into consideration(Logbook entry 450), and found out the beam is a little bit right on the mirror, left on the 10m target. So the dash line shows the situation now. So then we will move the PR and BS suspension mirror together to send the beam straight, after centered it on the window and the 10m target, we will try to center it on the 290m target.


At first we tried to move the beam in the horizontal direction out of dichroic mirror. But while we were moving, we thought that if the beam is elongated in the horizontal direction because of the mirror's wedge, when we scan the mirror maybe the beam shape will change. We scanned it,but the beam is too small,it was hard to say if it has some change or not, then we thought maybe we can put a lens with very short focal length on the BS bench to make the beam bigger in the short distance, but it was also hard to say since we scanned the beam in very small step, then on the BS bench the beam moved a lot and out of the lens, then we cannot exactly sure every time we put the lens in a good position to let the beam cross its center, if we did not cross center, the beam will also effected by the lens.
Then we followed our original plan to move the beam out of the dichroic in the horizontal direction, but it stopped by the mount of the mirror, so we tried in the vertical way, when we reached the pitch limitation of turning mirror, the beam was on the top edge of the dichroic, we thought it will be better if we move the beam totally outside the dichroic, so we decided to change a little bit the degree of last mirror on the bench. Before we changed the last mirror, we use the target to set the position of the green beam after the window for easier recover the beam. Then we let the beam go outside and far away from the BS chamber, to check the situation of it.
There were some strange situation that in the near field of the BS chamber, the beam elongation is much serious than the far field. We guessed that maybe because there are some overlap in the near field to make us feel the beam is longer, then in the far field, the other small beams separated from the main one. Actually, we did see that in the far field there were more small beams. But at least, the beam got better in this test.
The other situation I want to mention is that, I am not sure that the beam got better is only because of moving out from dichroic, we also have other changes in this test. Like I showed the simple drawing in the first picture, the circle 1,2,3 is the changed point.(It maybe a little bit hard to understand since I change the 3D situation into a 2D drawing, the red line is after we made the beam path goes up, so the red beam reached the higher position of the turning mirror and pass the dichroic from top of it)
1. The beam position on the window
2. The beam position on the turning mirror
3. The beam did not cross the dichroic
Also I showed the beam shape in two situation and in the near field and far field.
The second picture is I showed yesterday on the meeting for better understanding of the situation.

the He-Ne laser beam goes through the input mirror. Also in this case I centered the He-Ne laser beam
on the the first target in the tube and on the window at the entrance of the tube.
In order to do so the injection point from the back of the suspension is 21.8 cm in height and 2.4 cm
to the west.